OpenStudy (anonymous):

simplify without a negative exponent: 5^-2

OpenStudy (anonymous):

(1/5)^2

hartnn (hartnn):

\(\huge a^{-b}=\frac{1}{a^b}\) does this help ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1/25 is the final answer?

hartnn (hartnn):

yup :)

OpenStudy (anonymous):

thank you!

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

I have one more question..

hartnn (hartnn):

sure, ask.

OpenStudy (anonymous):

(y^2x^3)^4

hartnn (hartnn):

ok, use this \(\huge (ab)^n=a^n.b^n\)

hartnn (hartnn):

and then \(\huge (a^{b})^{n}=a^{nb}\)

hartnn (hartnn):

so whats \(\large (y^2x^3)^4\) get simplified to in 1st step ?

OpenStudy (anonymous):

(y^6x^7)

hartnn (hartnn):

nopes, u added the exponents.... according to 2nd property i have show, the exponents get multiplied....so can u try again ?

OpenStudy (callisto):

Sorry to interrupt a bit, but please start a new post for a new question next time. Thanks! Please continue with your discussion :)

OpenStudy (anonymous):

(y^8x^12)

hartnn (hartnn):

yup, thats correct now... and yes new question, new post, ok ?